OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3.

OpenStudy (anonymous):

Use point slope form y-y1=m(x-x1) just plug in your point (x1,y1) and the slope m and then simplify

hero (hero):

y = mx + b -5 = -3(2) + b -5 = -6 + b 6 - 5 = b 1 = b y =-3x + 1